OpenStudy (anonymous):

5/3=n/12 and 5/3 is negative solve each equation

OpenStudy (anonymous):

\[\frac{5}{3}=\frac{n}{12}\] 5*12=3n 60 = 3n 60/3=n n= 20

OpenStudy (anonymous):

each? there is only one. the method to solving this problem goes like so: first you need to multiply each side by the denominator of the other side. it should look like this: 12(-5/3)=(n/12)12 the 12's will cancel on the right giving you: -60/3=n now reduce the fraction and you get: -20=n

OpenStudy (anonymous):

yes n=-20

OpenStudy (anonymous):

yep its correct!!!